### v1

# Speeding up quantum measurement using space-time trade-off

Abstract
We present a scheme for speeding up quantum measurement. The scheme builds on previous protocols that entangle the system to be measured with ancillary systems. In the idealised situation of perfect entangling operations and no decoherence, it gives an exact space-time trade-off meaning the readout time reduces linearly with the number of ancilla. We verify this scheme is robust against experimental imperfections through numerical modelling of gate noise and readout errors, and under certain circumstances our scheme can even lead to better than linear improvement in the speed of measurement with the number of systems measured. This hardware-agnostic approach is broadly applicable to a range of quantum technology platforms and offers a route to accelerate mid-circuit measurement as required for effective quantum error correction.
I Introduction
High-fidelity readout of individual quantum systems is an essential requirement for emerging quantum technologies including precision metrology and quantum computing. Fast measurements are highly desirable, particularly in the context of mid-circuit measurements for quantum error correction where readout duration imposes a significant overhead [1, 2, 3, 4], however in practice there is a trade-off between measurement duration and quality. This places a lower bound on how fast a measurement can be performed to achieve a given fidelity. Even in cases where time is not a finite resource, such as the final state readout at the end of a computation, experimental imperfections and qubit dephasing in real systems impose an upper limit on readout time and hence measurement quality. In both scenarios, we seek a protocol able to simultaneously reduce readout duration and enhance the measurement fidelity.
There has been much previous work on how to improve the quality of measurements without simply increasing their duration. They can be split into two categories, passive methods, which correct errors on expectation values, [5, 6, 7, 8] and active methods, which correct errors on single-shot measurements [9, 10, 11, 12, 13, 14, 15]. For applications such as mid-circuit measurements, passive methods are insufficient as we need high quality results on a single-shot. While active methods have been shown to improve the quality of single-shot measurements, there has been little attention paid to how we can speed them up. The work towards faster measurements is a hard engineering challenge and thus far, progress has only been made for specific architectures [16, 17, 18, 19, 20, 21, 22, 23, 3, 12]. No thorough analysis has been performed at the abstract, hardware-agnostic level.
In this paper, we propose a scheme for speeding up quantum measurements. Our scheme builds on protocols in previous work [9, 10, 11, 12] where a repetition code is used to spread the information by entangling the system to be measured with ancillary qubits. The key novelty in our work is to harness these protocols to provide a scheme for reducing measurement time while maintaining (or even improving) measurement quality. While the simplest embodiment of our scheme describes using multiple measurement devices, in fact our scheme can also be implemented using only a single measurement device coupled to the target and ancillary qubits. A further impact of reducing measurement time is to suppress decoherence processes, which allows us to outperform the maximum quality imposed by the upper bound on measurement time. The effect of the scheme is to trade-off space (number of ancillary systems) against the time taken to do the measurement.
A vital aspect of the scheme is that the trade-off is robust against imperfections in its implementation. We illustrate this by detailed numerical simulations of a natural, simple but characteristic model of noise and decoherence. Crucially, we note that in certain circumstances with realistic noise parameters our scheme can lead to a better than linear improvement in the speed of measurement with the number of systems measured, achieving optimal readout parameters with a modest number of additional qubits.
In summary, our scheme is hardware agnostic and provides a robust, powerful and versatile method for speeding up quantum measurements using ancillary systems.
II Case study: Poisson measurement statistics
In order to present the main idea, we will first focus on state dependent fluorescence qubit measurement which is the primary method used in trapped ion [2, 24] and neutral atom [1, 25, 26, 27] quantum computing.
Qubit information is encoded in either a “dark” or “bright” state of an ion or atom. When illuminated by a detection laser, the atom will be projected into one of these states and will then scatter photons. A measurement device (detector) will count the number $k$ of photons detected. We have in mind to count photons for a time $t$. The bright state will scatter many photons with a mean of $\nu_{1}(t)$ while the dark state will scatter relatively fewer photons with a mean of $\nu_{0}(t)$ where $\nu_{1}(t)\geq\nu_{0}(t)$. To discriminate the bright and dark states we define a threshold value $\eta(t)$ where $\nu_{0}(t)<\eta(t)<\nu_{1}(t)$. If the number of photons counted is $k<\eta(t)$ then we assign the outcome 00 and if $k\geq\eta(t)$ we assign the outcome $1$. From now on we refer to the bright state as |$\ket{1}$ and the dark state as |$\ket{0}$.
Let us denote by $\mathcal{P}_{\ket{j},t}^{(n)}(k)$ the probability distribution of the random variable $K_{\ket{j},t}^{(n)}$, defining to the number of photons counted, when we use $n$ detectors for time $t$ and the input state is |$\ket{j}$, $j\in\{0,1\}$. Initially we consider $n=1$.
In an idealised situation, the number of photons counted will follow a Poisson distribution
$\mathcal{P}_{\ket{j},t}^{(1)}(k)=L_{\nu_{j}(t)}(k),$ (1)
where $L_{\omega}(k)=\frac{\omega^{k}}{k!}e^{-\omega}$. Recall the mean and variance of a Poisson distribution coincide and are both equal, $\langle K\rangle_{L_{\omega}}=\Delta K^{2}_{L_{\omega}}=\omega$. The two distributions, $\mathcal{P}_{\ket{0},t}^{(1)}$ and $\mathcal{P}_{\ket{1},t}^{(1)}$, will have an overlap which is the source of error in measurement quality when assigning outcomes; no threshold value can be defined which will give you a perfect measurement by correctly discriminating the two states.
A key property of the Poisson distribution is that the parameter defining it is additive. If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution,
$L_{\nu_{j}(t_{1}+t_{2})}(k)$ $=\sum_{k=k_{1}+k_{2}}L_{\nu_{j}(t_{1})}(k_{1})L_{\nu_{j}(t_{2})}(k_{2}),$
$=(L_{\nu_{j}(t_{1})}*L_{\nu_{j}(t_{2})})(k),$
$=L_{\nu_{j}(t_{1})+\nu_{j}(t_{2})}(k),$
i.e. the Poisson distribution convolves to itself. This implies that $\nu_{j}(t)=\mu_{j}t$, for a rate $\mu_{j}$ independent of $t$. In particular the convolution property means that $\nu_{j}(2t)=2\nu_{j}(t)$.
Although one is generally interested in how good a measurement is for an arbitrary state, a commonly used figure of merit is the distinguishability, as quantified by the signal-to-noise-ratio (SNR), of the |$\ket{0}$ and |$\ket{1}$ state probability distributions. For two arbitrary probability distributions $P_{0}(x)$ and $P_{1}(x)$, the SNR is
$S(P_{0},P_{1})=\frac{2|\langle X\rangle_{P_{0}}-\langle X\rangle_{P_{1}}|}{\Delta X_{P_{0}}+\Delta X_{P_{1}}}.$ (2)
The larger the SNR, the better the distinguishability and the better the measurement quality.
In our case, for the two Poisson distributions in (1) we find
$S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})=\frac{2\sqrt{t}|\mu_{0}-\mu_{1}|}{\sqrt{\mu_{0}}+\sqrt{\mu_{1}}}.$ (3)
Thus the quality of the measurement is proportional to $\sqrt{t}$ and therefore increases with measurement duration. This provides a quantification of the inverse trade-off between the speed and quality of a measurement.
II.1 Exact space-time trade-off in the noiseless case
We now introduce a scheme which, rather than trading time for quality as above, allows us to trade time for “space” – by which we mean number of qubits. In particular, this provides a means to realise high-quality fast measurements, at the cost of using additional ancillas. Our scheme is outlined in Fig. 1, and employs a simple ‘cloning’ protocol. We apply a CNOT gate between our target qubit and an ancillary qubit initialised in the state |$\ket{0}$. We assume CNOTs can be performed in a negligibly-short time compared to the measurement duration. We measure the ancilla and the original qubit then assign the result as either 00 or $1$, representing the outcome of the measurement on the target qubit. In the ideal case, when measured for a time $t$, we will now show that this scheme gives identical measurement statistics to that of a standard measurement on a single qubit measured for a time $2t$. This thus effectively halves the measurement time.
Figure 1: Space-time trade-off scheme. Example using one additional ancillary qubit. We entangle the input qubit with an ancillary qubit in state |$\ket{0}$ using a CNOT gate. We then measure each qubit for a time $t$, obtaining results $k_{1}$ and $k_{2}$ respectively. The statistics of the sum $k=k_{1}+k_{2}$ coincides with the statistics for a single measurement of time $2t$. We thus effectively speed up the measurement by a factor of 2 using this scheme.
For now, for clarity of presentation, we analyse the case of independent detectors for each qubit, but come back to this point below. We denote by $k_{1}$ the number of photons detected by the 1st detector, and $k_{2}$ for the second. The measurement statistics for the pair of outcomes $(k_{1},k_{2})$ are
|$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2}),$ (4)
where |$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})$ is the measurement statistics for the input state |$\ket{jj}$ when each measurement lasts for a time $t$. The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, hence
$\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k).$ (5)
An important feature is present here; since the probability distribution only considers the total sum of emissions from both fluorescing qubits, we could instead consider the distribution as describing the number of photons counted from a single detector which is collecting all the emitted photons, rather than collecting separately and then adding.
The key insight is that, due to the convolution property of the Poisson, measuring a single qubit for time $2t$ is equivalent to summing the results from two measurements, each of time $t$, on two entangled qubits. Specifically
$\mathcal{P}_{\ket{j},t}^{(2)}(k)=\mathcal{P}_{\ket{j},2t}^{(1)}(k).$ (6)
This result shows that we can speed the measurement up by a factor of $2$, by performing a statistically identical measurement in time $\frac{t}{2}$ as was previously performed in time $t$. Since the measurements are identical, the quality, as quantified by the SNR (or any other figure of merit), will be constant while the measurement is performed faster.
This result generalises to $N$ qubits where we entangle $N-1$ ancillary qubits, in the initial state |$\ket{0}$, to the target qubit using CNOT gates. In this case we find
$\mathcal{P}_{\ket{j},t}^{(N)}(k)=(L_{\mu_{j}t}^{*N})(k)=L_{N\mu_{j}t}(k)=\mathcal{P}_{\ket{j},Nt}^{(1)}(k),$ (7)
where $k$ is the total sum of photons counted and $(f^{*N})(w)$ is the $N$-fold convolution of the function $f(v)$. This shows that we can keep the quality of a measurement constant, and speed it up arbitrarily, from time $t$ to time $t/N$, by making use of $N$ measurements in parallel.
A useful ratio for comparison with more realistic devices later on is $t^{(1)}_{S}/t^{(N)}_{S}$ where $t^{(N)}_{S}$ is the time it takes an $N$ qubit scheme to reach an SNR of $S$. In this case we have
$t^{(1)}_{S}/t^{(N)}_{S}=N$ (8)
for all $S$, so the improvement in speed is exactly linear in qubit number. The speed up from our scheme is shown in Fig. 2 (a) and (b). The parameters used for the plots are experimentally relevant for neutral atom quantum computers and are found in [28].
III General space-time trade-off
A key insight is that the above analysis is much more general than just Poisson statistics arising from counting photons. In fact, the above scheme will lead to a space-time trade-off in a wide range of situations, with very few assumptions about the characteristics of the measuring device.
We will assume that we have a measurement procedure that is time dependent, meaning that the duration of the measurement can be changed, which will change the quality of the measurement. Let us assume that the statistics of this measurement (with outcome $x$) when measuring the two basis states for a time $t$ is given by $\mathcal{P}_{\ket{j},t}^{(1)}(x)$, where the superscript reminds us that this is without using the measurement scheme, i.e. we are performing a single measurement. For this measurement we can calculate the SNR (2) between these two distributions. Let us assume the dependence with time, to leading order, is
$S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}.$ (9)
For the case of Poisson statistics above, we saw that $\alpha~{}=~{}\frac{1}{2}$, however we may imagine an alternative dependence on time.
Figure 2: Space-time trade-off for idealised and realistic devices. Plots show how faster measurements can be made without sacrificing distinguishability by introducing more qubits. The parameters used for the plots are experimentally relevant and are found in [28]. In plots a) and b), we use Poisson distributed measurement statistics (7) with $\mu_{0}=3.5\text{ms}^{-1}$ and $\mu_{1}=14\text{ms}^{-1}$ for different numbers of qubits $N$. a) shows the SNR (3) against time from $N=1$ to $N=5$ qubits. The SNR will always increase with time, however the rate at which the SNR improves increases with qubit number. The dashed grey lines are lines of constant SNR at $S=8,9.5,10.6$ and are plotted in b) which shows the time of measurement $t$ against number of qubits from $N=1$ to $N=10$ qubits. The inset shows the same points but for the inverse time $1/t$. The inset clearly shows the linear relationship implied by (8). In plots c) and d) we show similar results but for realistic devices. In this case, a higher maximum distinguishability can be achieved by introducing more qubits and faster measurements can be made while also improving distinguishability. These plots are generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1), found in the Appendix, with the cascade compilation of noisy CNOT gates defined in the Supplementary material. We use input parameters $\mu_{0}=3.5\text{ms}^{-1}$, $\mu_{1}=14\text{ms}^{-1}$, $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate. c) shows the SNR (2) against time for different numbers of qubits $N$. The SNR increases with time up to a maximum value then decreases due to decoherence. The rate at which the SNR improves increases with qubit number, as does the maximum SNR achieved. The dashed grey lines are lines of constant SNR at $S=8,9.5,10.6$ and are plotted in d) which shows the time of measurement $t$, when the defined SNR is first achieved, against number of qubits $N$. Notice how $S=9.5$ has no point for $N=1$ and $S=10.6$ has no points for $N=1,2$; this means these values cannot be reached without more qubits. The inset shows the same points plotted for inverse time $1/t$. The points do not follow a straight line so the space-time trade-off is not linear with qubit number $N$. The Appendix includes figures with extra details for c) and d). Fig. 3 demonstrates the increase in the peak SNR value with qubit number and Fig. 4 demonstrates the increase of the speed up with qubit number.
Now, in the above, we saw that in an idealised setting, involving no noise in the scheme, that by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (7). Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, we nevertheless have the fundamental property that the expectation and variance add,
$\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}},$ $\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}.$ (10)
It therefore follows that the SNR for this measurement scheme will be
$S(\mathcal{P}_{\ket{0},t}^{(N)},\mathcal{P}_{\ket{1},t}^{(N)})=\sqrt{N}S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}\sqrt{N}.$ (11)
Thus, as long as $t^{\alpha}\sqrt{N}$ is constant, then the quality of the measurement is constant, and hence we still obtain a space-time trade-off. In particular, if $t_{S}^{(1)}$ is the duration of the measurement which achieves an SNR of $S$, then using the scheme, we can speed this up to a measurement time of $t^{(N)}_{S}$ given by
$\frac{t_{S}^{(1)}}{t^{(N)}_{S}}=N^{1/2\alpha}.$ (12)
This recovers (8) in the case $\alpha=1/2$ but shows more generally, if $\alpha>1/2$, we will still achieve a trade-off between space and time using this scheme, albeit not as good as (8). However, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (8).
One application of this general result is that it allows us to take into account decoherence in the measurement process. In the next section we address the important question of how imperfections in the scheme affect its overall performance.
IV Space-time trade-off for more realistic devices
For our scheme to be useful in practical applications it is crucial to show that the speed-up is robust to realistic imperfection. To this end, we now describe a simple noise model to take into account these dominant sources of error. Since ancilla qubits are initialised in state |$\ket{0}$, imperfections in the CNOT gate operations can be modelled by considering errors on state |$\ket{10}$ being mapped to state |$\ket{00}$ with error probability p (see eg. [29]). Additionally, we consider imperfections in qubit readout caused by qubits in state |$\ket{1}$ falling into |$\ket{0}$ during the measurement, such as due to depuming by off-resonant scattering in atomic fluorescence measurements [30]. This error is included by applying a probabilistic decay at constant rate $\lambda$ during the measurement. See the Appendix for further details on our noise model.
The effects are summarised in Figures 2c and 2d, in direct comparison to the perfect case. The SNR is calculated using (29) in the Supplementary material, which gives the SNR of a realistic $N$ qubit scheme entirely in terms of the expectation and variance of the single qubit states |$\ket{0}$ and |$\ket{1}$ and the expectation and variance of the CNOT noise.
We see that we still obtain a significant space-time trade-off, allowing for significantly faster measurements, even in the presence of noise, demonstrating the robustness of this scheme.
There are other notable points worth highlighting. First, unlike in the perfect case, the SNR does not increase indefinitely. It reaches a peak before starting to decrease, due to decoherence. We see that the peak value is increased by schemes using more qubits. Thus not only does the scheme speed up the measurement while keeping the quality constant, it can also improve the measurement at the same time. This is shown in Fig. 3 in the Appendix in more detail. Similar to the perfect case, the rate of SNR increase improves with $N$, but it can now be a “super-improvement”; for small enough values of $N$ and CNOT noise $p$, the rate of speed up is better than (8) for the relevant time scales, that is the times before decoherence causes the measurement to deteriorate, as demonstrated by Fig. 4 in the Appendix.
V Conclusion
In this paper we have presented a general scheme for speeding up quantum measurement while maintaining quality. This is achieved by means of a space-time trade-off, using additional ancillary qubits in place of longer measurement time. We presented the main idea using ideal Poisson measurement statistics but emphasise that the scheme is much more general than this. In particular, by using characteristic noise models for the CNOTs and decoherence, the scheme was shown to be robust to imperfections while simultaneously able to achieve a “super-improvement” in speed-up, since it acts to suppresses decoherence, allowing for a higher achievable maximum quality. While we have illustrated the scheme with systems of qubits, it should be clear from the discussions above that the scheme is equally applicable to higher dimensional systems.
Since speeding up quantum measurements is a highly relevant problem, we envisage that our scheme will have significant applications across quantum technologies.
VI Note added
During preparation of this manuscript we became aware of the independent work of Petrosyan, Norrell, Poole and Saffman [31] which contains related ideas.
Acknowledgements.
VII Acknowledgements
We acknowledge support from the EPSRC through the following grants: EP/T005386/1, EP/T001062/1, EP/Y005058/2, EP/Y004655/1, EP/L015730/1 and EP/R043957/1. We also acknowledge support from Agence Nationale de la Recherche (ANR-22-PNCQ-0002), the Royal Society (URF NFQI) and from a CIFAR Azrieli Global Scholarship.
VIII Appendices
VIII.1 Appendix on noise model
We consider a noise model with two key ingredients: (i) Noise in the CNOT gates applied; (ii) decoherence of the individual systems. We outline here these two elements separately.
For the CNOT noise, we consider a simple model inspired by neutral atom quantum computers. In this model, whenever the system is in state |$\ket{1}$ and a CNOT is applied, there is a chance of failure, which causes the system to decay to |$\ket{0}$. More formally, we define a probability $p$, that immediately before a CNOT is implemented, |$\ket{1}$ will be sent to |$\ket{0}$. If this happens, the CNOT can not flip the state |$\ket{0}$ back to |$\ket{1}$. The noisy CNOT channel $\widetilde{C_{p}}$ acts on the basis states, where the first qubit is the control qubit, as
$\widetilde{C_{p}}(\ket{00}\bra{00})$ =|$=\ket{00}\bra{00}$ |
$\widetilde{C_{p}}(\ket{10}\bra{10})$ $=p\ket{00}\bra{00}+(1-p)\ket{11}\bra{11}.$ | . (13)
Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{0}$ and |$\ket{1}$ of the system qubit. Furthermore, we assume our ancillary qubits are initialised perfectly in |$\ket{0}$ so we are only interested in how the noisy CNOT gates will act on the states |$\ket{00}$ and |$\ket{10}$. Thus (VIII.1) defines the action on all states that occur in the scheme.
When the system is in the state |$\ket{0}$, and we perform $N$ measurements, the initial state will be |$\ket{0}^{\otimes N}$. In this case, the action of $\widetilde{C_{p}}$ is to leave the state of all $N$ systems as |$\ket{0}^{\otimes N}$. When the system is in the state |$\ket{1}$, the initial state is |$\ket{1}\ket{0}^{\otimes(N-1)}$. Our scheme only depends upon the total number of photons emitted from each qubit (and does not depend upon which qubit emitted which photon). We thus only need to calculate the probability distribution for the total number of qubits in state |$\ket{1}$ after applying the $\widetilde{C_{p}}$ gates.
There are numerous ways the gates can be compiled and different compilations will give different probability distributions for the number of qubits in state |$\ket{1}$. These different compilations are explored in more detail in the supplementary material. Let us denote by $\mathcal{T}^{(N)}_{p}(q)$ the probability distribution of the random variable $Q^{(N)}_{p}$ defining the number of qubits in state |$\ket{1}$ after applying the noisy CNOT gates. Any qubits not in state |$\ket{1}$ are in state |$\ket{0}$. Combining this with the probability distribution for $N$ qubits emitting $k$ photons (7), we find the resultant distribution using our scheme with noisy CNOT gates to be
$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k)$
$\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(L_{\mu_{1}t}^{*q}*L_{\mu_{0}t}^{*N-q})(k)$
$=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)L_{(q\mu_{1}+(N-q)\mu_{0})t}(k).$ (14)
For the decoherence noise, we consider a model whereby qubits in state |$\ket{1}$ probabilistically decay at a constant rate $\lambda$ to state |$\ket{0}$. After this happens, the qubit will emit photons at a different rate so the resultant distribution of photons will not be the same as a perfect Poisson. We are interested in the probability distribution for the number of photons emitted $k$ in time $t$. We model this behaviour by finding the probability a qubit in state |$\ket{1}$ will emit $k^{\prime}$ photons in time $t^{\prime}$ before decaying to |$\ket{0}$ and then emitting $k-k^{\prime}$ photons in time $t-t^{\prime}$, then integrating over $t^{\prime}$ and summing over $k^{\prime}$. We find the resultant probability distribution to be
[…]
The final integral can be readily evaluated numerically.
Figure 3: Increase in peak SNR from Fig. 2 (c) and (d). These Plots are generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1) with the cascade compilation of noisy CNOT gates defined in the Supplementary material. We use input parameters $\mu_{0}=3.5$, $\mu_{1}=14$, $\lambda=0.0041$ and $p=0.01$. We plot the peak SNR $S_{max}$ for $N=1$ to $N=10$ qubits. While more qubits result in a higher peak SNR, this trend flattens out for higher qubit numbers.
Figure 4: Speed up from Fig. 2 (c) and (d). Plot generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1) with the cascade compilation of noisy CNOT gates defined in the Supplementary material. We use input parameters $\mu_{0}=3.5$, $\mu_{1}=14$ and $\lambda=0.0041$. We plot the ratio $t^{(1)}_{S}/t^{(N)}_{S}$ for two different values of $p$, at $S=8$, for $N=1$ to $N=10$ qubits and the linear relationship $t^{(1)}_{S}/t^{(N)}_{S}=N$ (lime dashed line) achieved by the perfect Poisson distribution (8). The value $p=0.01$ (blue dots) is used in Fig. 2 (c) and (d). The points are increasing, which shows we get a speed up, however they are below the dashed line so the speed-up is worse than the linear case. On the other hand, the values for $p=0.001$ (red crosses) are above the dashed line which demonstrates that, for CNOTs with a low enough failure rate, we can beat the decoherence of the qubit; it is possible to achieve a “super-improvement”.
Our complete noise model consists of combining the noisy CNOT protocol (VIII.1) with the decaying Poisson distribution (15). Altogether, we we thus find that the probability distributions for the number of photons, using our scheme with $N$ systems measured, is given by
$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k)$
$\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(\mathcal{W}_{\mu_{1},\mu_{0},\lambda,t}^{*q}*L_{(N-q)\mu_{0}t})(k).$ (16)
These are used to produce the results in Fig. 2 (c) and (d).

### v2

@@
 The scheme builds on previous protocols that entangle the system to be measured with ancillary systems.
-In the idealised situation of perfect entangling operations and no decoherence, it gives an exact space-time trade-off meaning the readout time reduces linearly with the number of ancilla.
+In the idealised situation of perfect entangling operations and no decoherence, it gives an exact space-time trade-off meaning the readout speed increases linearly with the number of ancilla.
 We verify this scheme is robust against experimental imperfections through numerical modelling of gate noise and readout errors, and under certain circumstances our scheme can even lead to better than linear improvement in the speed of measurement with the number of systems measured.
@@
 II Case study: Poisson measurement statistics
-In order to present the main idea, we will first focus on state dependent fluorescence qubit measurement which is the primary method used in trapped ion [2, 24] and neutral atom [1, 25, 26, 27] quantum computing.
+In order to present the main idea, we will first focus on state dependent fluorescence qubit measurement which is the primary method used in trapped-ion [2, 24] and neutral-atom [1, 25, 26, 27] quantum computing.
 Qubit information is encoded in either a “dark” or “bright” state of an ion or atom.
@@
 From now on we refer to the bright state as |$\ket{1}$ and the dark state as |$\ket{0}$.
-Let us denote by $\mathcal{P}_{\ket{j},t}^{(n)}(k)$ the probability distribution of the random variable $K_{\ket{j},t}^{(n)}$, defining to the number of photons counted, when we use $n$ detectors for time $t$ and the input state is |$\ket{j}$, $j\in\{0,1\}$.
+In all that follows, we denote $j\in\{0,1\}$.
+Let us denote by $\mathcal{P}_{\ket{j},t}^{(n)}(k)$ the probability distribution of the random variable $K_{\ket{j},t}^{(n)}$, defining the number of photons counted, when we use $n$ detectors for time $t$ and the input state is |$\ket{j}$.
 Initially we consider $n=1$.
@@
 The two distributions, $\mathcal{P}_{\ket{0},t}^{(1)}$ and $\mathcal{P}_{\ket{1},t}^{(1)}$, will have an overlap which is the source of error in measurement quality when assigning outcomes; no threshold value can be defined which will give you a perfect measurement by correctly discriminating the two states.
-A key property of the Poisson distribution is that the parameter defining it is additive.
 If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution,
@@
 In particular the convolution property means that $\nu_{j}(2t)=2\nu_{j}(t)$.
-Although one is generally interested in how good a measurement is for an arbitrary state, a commonly used figure of merit is the distinguishability, as quantified by the signal-to-noise-ratio (SNR), of the |$\ket{0}$ and |$\ket{1}$ state probability distributions.
-For two arbitrary probability distributions $P_{0}(x)$ and $P_{1}(x)$, the SNR is
-$S(P_{0},P_{1})=\frac{2|\langle X\rangle_{P_{0}}-\langle X\rangle_{P_{1}}|}{\Delta X_{P_{0}}+\Delta X_{P_{1}}}.$ (2)
+Although one is generally interested in how good a measurement is for an arbitrary state, a commonly used figure of merit is the distinguishability, as quantified by the signal-to-noise-ratio (SNR), of the ${\mathcal{P}_{\ket{0},t}^{(n)}(k)}$ and ${\mathcal{P}_{\ket{1},t}^{(n)}(k)}$ probability distributions.
+The SNR is
+$S\left(\mathcal{P}_{\ket{0},t}^{(n)},\mathcal{P}_{\ket{1},t}^{(n)}\right)=\frac{2\left|\langle X\rangle_{\mathcal{P}_{\ket{0},t}^{(n)}}-\langle X\rangle_{\mathcal{P}_{\ket{1},t}^{(n)}}\right|}{\Delta X_{\mathcal{P}_{\ket{0},t}^{(n)}}+\Delta X_{\mathcal{P}_{\ket{1},t}^{(n)}}}.$ (2)
 The larger the SNR, the better the distinguishability and the better the measurement quality.
-In our case, for the two Poisson distributions in (1) we find
-$S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})=\frac{2\sqrt{t}|\mu_{0}-\mu_{1}|}{\sqrt{\mu_{0}}+\sqrt{\mu_{1}}}.$ (3)
+In our case, for the two Poisson distributions, we find
+$S\left(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)}\right)=\frac{2\sqrt{t}|\mu_{0}-\mu_{1}|}{\sqrt{\mu_{0}}+\sqrt{\mu_{1}}}.$ (3)
 Thus the quality of the measurement is proportional to $\sqrt{t}$ and therefore increases with measurement duration.
 This provides a quantification of the inverse trade-off between the speed and quality of a measurement.
-II.1 Exact space-time trade-off in the noiseless case
+III Exact space-time trade-off in the idealised Poisson case
 We now introduce a scheme which, rather than trading time for quality as above, allows us to trade time for “space” – by which we mean number of qubits.
 In particular, this provides a means to realise high-quality fast measurements, at the cost of using additional ancillas.
-Our scheme is outlined in Fig.
-1, and employs a simple ‘cloning’ protocol.
+See Fig.
+1.
 We apply a CNOT gate between our target qubit and an ancillary qubit initialised in the state |$\ket{0}$.
-We assume CNOTs can be performed in a negligibly-short time compared to the measurement duration.
+In this section we consider the idealised situation that the CNOTs can be performed perfectly and in a negligibly-short time compared to the measurement duration.
 We measure the ancilla and the original qubit then assign the result as either 00 or $1$, representing the outcome of the measurement on the target qubit.
 In the ideal case, when measured for a time $t$, we will now show that this scheme gives identical measurement statistics to that of a standard measurement on a single qubit measured for a time $2t$.
-This thus effectively halves the measurement time.
+The measurement time is effectively halved.
 Figure 1: Space-time trade-off scheme.
-Example using one additional ancillary qubit.
+Idealised version using one additional ancillary qubit.
 We entangle the input qubit with an ancillary qubit in state |$\ket{0}$ using a CNOT gate.
@@
 The statistics of the sum $k=k_{1}+k_{2}$ coincides with the statistics for a single measurement of time $2t$.
-We thus effectively speed up the measurement by a factor of 2 using this scheme.
+We thus speed up the measurement by a factor of 2.
 For now, for clarity of presentation, we analyse the case of independent detectors for each qubit, but come back to this point below.
-We denote by $k_{1}$ the number of photons detected by the 1st detector, and $k_{2}$ for the second.
-The measurement statistics for the pair of outcomes $(k_{1},k_{2})$ are
-|$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2}),$ (4)
-where |$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})$ is the measurement statistics for the input state |$\ket{jj}$ when each measurement lasts for a time $t$.
-The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, hence
-$\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k).$ (5)
-An important feature is present here; since the probability distribution only considers the total sum of emissions from both fluorescing qubits, we could instead consider the distribution as describing the number of photons counted from a single detector which is collecting all the emitted photons, rather than collecting separately and then adding.
-The key insight is that, due to the convolution property of the Poisson, measuring a single qubit for time $2t$ is equivalent to summing the results from two measurements, each of time $t$, on two entangled qubits.
-Specifically
-$\mathcal{P}_{\ket{j},t}^{(2)}(k)=\mathcal{P}_{\ket{j},2t}^{(1)}(k).$ (6)
-This result shows that we can speed the measurement up by a factor of $2$, by performing a statistically identical measurement in time $\frac{t}{2}$ as was previously performed in time $t$.
+We denote by $k_{1}$ and $k_{2}$ the number of photons detected from the first and second qubit detectors respectively and we input the state |$\ket{\psi}=\ket{j}$ into our scheme.
+The measurement statistics for the pair of outcomes $(k_{1},k_{2})$, given we are measuring the post-CNOT state |$\ket{jj}$ for a time $t$, are
+|$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2}).$ (4)
+The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, thus ${\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k)}$.
+An important feature is present here; since the probability distribution only considers the total sum of emissions from both fluorescing qubits, we can consider the distribution as describing the number of photons counted from a single detector which is collecting all the emitted photons, rather than collecting separately and then adding.
+The key insight is that, due to the convolution property of the Poisson, measuring a single qubit for time $2t$ is equivalent to summing the results from two measurements, of time $t$, on two entangled qubits.
+Specifically ${\mathcal{P}_{\ket{j},t}^{(2)}(k)=\mathcal{P}_{\ket{j},2t}^{(1)}(k).}$ This result shows that we can speed up the measurement, by a factor of $2$, by performing a statistically identical measurement in time $\frac{t}{2}$ as was previously performed in time $t$.
 Since the measurements are identical, the quality, as quantified by the SNR (or any other figure of merit), will be constant while the measurement is performed faster.
@@
 $\mathcal{P}_{\ket{j},t}^{(N)}(k)=(L_{\mu_{j}t}^{*N})(k)=L_{N\mu_{j}t}(k)=\mathcal{P}_{\ket{j},Nt}^{(1)}(k),$ (5)
-where $k$ is the total sum of photons counted and $(f^{*N})(w)$ is the $N$-fold convolution of the function $f(v)$.
+where $k$ is the total number of photons counted and $(f^{*N})(w)$ is the $N$-fold convolution of the distribution $f(v)$.
 This shows that we can keep the quality of a measurement constant, and speed it up arbitrarily, from time $t$ to time $t/N$, by making use of $N$ measurements in parallel.
-A useful ratio for comparison with more realistic devices later on is $t^{(1)}_{S}/t^{(N)}_{S}$ where $t^{(N)}_{S}$ is the time it takes an $N$ qubit scheme to reach an SNR of $S$.
-In this case we have
+A useful ratio for comparison with more realistic devices later on is $t^{(1)}_{S}/t^{(N)}_{S}$ where $t^{(N)}_{S}$ is the time it takes an $N$ qubit scheme to reach an SNR of $S$. In this case
 $t^{(1)}_{S}/t^{(N)}_{S}=N$ (6)
-for all $S$, so the improvement in speed is exactly linear in qubit number.
-The speed up from our scheme is shown in Fig.
-2 (a) and (b).
-The parameters used for the plots are experimentally relevant for neutral atom quantum computers and are found in [28].
-III General space-time trade-off
-A key insight is that the above analysis is much more general than just Poisson statistics arising from counting photons.
-In fact, the above scheme will lead to a space-time trade-off in a wide range of situations, with very few assumptions about the characteristics of the measuring device.
+for all $S$, so the improvement in speed is exactly linear in qubit number. The speed up from our scheme is shown in Figs. 2a and 2b.
+IV General space-time trade-off
+A key insight is that the scheme above applies much more generally than just to Poisson statistics arising from counting photons.
+In fact, the scheme gives a space-time trade-off in a wide range of measuring devices.
 We will assume that we have a measurement procedure that is time dependent, meaning that the duration of the measurement can be changed, which will change the quality of the measurement.
@@
 $S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}.$ (7)
-For the case of Poisson statistics above, we saw that $\alpha~{}=~{}\frac{1}{2}$, however we may imagine an alternative dependence on time.
+For the case of Poisson statistics above we saw that $\alpha~{}=~{}1/2$ however we may imagine an alternative dependence on time.
 Figure 2: Space-time trade-off for idealised and realistic devices.
-Plots show how faster measurements can be made without sacrificing distinguishability by introducing more qubits.
-The parameters used for the plots are experimentally relevant and are found in [28].
+The parameters used for the plots are experimentally relevant and are taken from [28].
 In plots a) and b), we use Poisson distributed measurement statistics (5) with $\mu_{0}=3.5\text{ms}^{-1}$ and $\mu_{1}=14\text{ms}^{-1}$ for different numbers of qubits $N$.
-a) shows the SNR (3) against time from $N=1$ to $N=5$ qubits.
-The SNR will always increase with time, however the rate at which the SNR improves increases with qubit number.
-The dashed grey lines are lines of constant SNR at $S=8,9.5,10.6$ and are plotted in b) which shows the time of measurement $t$ against number of qubits from $N=1$ to $N=10$ qubits.
-The inset shows the same points but for the inverse time $1/t$.
-The inset clearly shows the linear relationship implied by (8).
-In plots c) and d) we show similar results but for realistic devices.
+a) SNR (2) against time from $N=1$ to $N=5$ qubits.
+The SNR always increases with time and qubit number and continues the trend beyond what is shown on the graph.
+b) time of measurement $t$ against number of qubits from $N=1$ to $N=10$ qubits for the SNR values on the grey dashed lines in a).
+Inset: same points but for the inverse time $1/t$, demonstrating the linear relationship (6).
+c) and d) show similar results but for realistic devices.
 In this case, a higher maximum distinguishability can be achieved by introducing more qubits and faster measurements can be made while also improving distinguishability.
-These plots are generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1), found in the Appendix, with the cascade compilation of noisy CNOT gates defined in the Supplementary material.
-We use input parameters $\mu_{0}=3.5\text{ms}^{-1}$, $\mu_{1}=14\text{ms}^{-1}$, $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate.
-c) shows the SNR (2) against time for different numbers of qubits $N$.
-The SNR increases with time up to a maximum value then decreases due to decoherence.
-The rate at which the SNR improves increases with qubit number, as does the maximum SNR achieved.
-The dashed grey lines are lines of constant SNR at $S=8,9.5,10.6$ and are plotted in d) which shows the time of measurement $t$, when the defined SNR is first achieved, against number of qubits $N$.
-Notice how $S=9.5$ has no point for $N=1$ and $S=10.6$ has no points for $N=1,2$; this means these values cannot be reached without more qubits.
-The inset shows the same points plotted for inverse time $1/t$.
-The points do not follow a straight line so the space-time trade-off is not linear with qubit number $N$.
-The Appendix includes figures with extra details for c) and d).
-Fig.
-3 demonstrates the increase in the peak SNR value with qubit number and Fig.
-4 demonstrates the increase of the speed up with qubit number.
-Now, in the above, we saw that in an idealised setting, involving no noise in the scheme, that by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (7).
-Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, we nevertheless have the fundamental property that the expectation and variance add,
-$\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}},$ $\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}.$ (10)
-It therefore follows that the SNR for this measurement scheme will be
-$S(\mathcal{P}_{\ket{0},t}^{(N)},\mathcal{P}_{\ket{1},t}^{(N)})=\sqrt{N}S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}\sqrt{N}.$ (11)
+These plots are generated using the noise and decoherence model (IX), found in the Appendix.
+We use input parameters $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate.
+In c), the SNR increases with time up to a maximum value then decreases due to decoherence.
+The rate at which the SNR improves increases with qubit number, as does the peak SNR achieved.
+Inset: points in the inset no longer follow a straight line, indicating that the space-time trade-off is not linear with qubit number $N$.
+In the above we saw that, in an idealised version of the scheme involving perfect CNOTs, by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (5).
+Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, for perfect CNOTs we nevertheless have the fundamental property that the expectation and (due to independence) variance add,
+$\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}},$ $\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}.$
+It therefore follows that the SNR, using the scheme, will be
+$S\left(\mathcal{P}_{\ket{0},t}^{(N)},\mathcal{P}_{\ket{1},t}^{(N)}\right)=\sqrt{N}S\left(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)}\right)\sim t^{\alpha}\sqrt{N}.$ (8)
 Thus, as long as $t^{\alpha}\sqrt{N}$ is constant, then the quality of the measurement is constant, and hence we still obtain a space-time trade-off.
 In particular, if $t_{S}^{(1)}$ is the duration of the measurement which achieves an SNR of $S$, then using the scheme, we can speed this up to a measurement time of $t^{(N)}_{S}$ given by
-$\frac{t_{S}^{(1)}}{t^{(N)}_{S}}=N^{1/2\alpha}.$ (12)
+$t_{S}^{(1)}/t^{(N)}_{S}=N^{1/(2\alpha)}.$ (9)
 This recovers (6) in the case $\alpha=1/2$ but shows more generally, if $\alpha>1/2$, we will still achieve a trade-off between space and time using this scheme, albeit not as good as (6).
-However, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (8).
-One application of this general result is that it allows us to take into account decoherence in the measurement process.
-In the next section we address the important question of how imperfections in the scheme affect its overall performance.
-IV Space-time trade-off for more realistic devices
+Nevertheless, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (6).
+V Space-time trade-off in more realistic settings
 For our scheme to be useful in practical applications it is crucial to show that the speed-up is robust to realistic imperfection.
-To this end, we now describe a simple noise model to take into account these dominant sources of error.
+To this end, we now describe a simple noise model to take into account the dominant sources of error.
 Since ancilla qubits are initialised in state |$\ket{0}$, imperfections in the CNOT gate operations can be modelled by considering errors on state |$\ket{10}$ being mapped to state |$\ket{00}$ with error probability p (see eg.
@@
 See the Appendix for further details on our noise model.
-The effects are summarised in Figures 2c and 2d, in direct comparison to the perfect case.
-The SNR is calculated using (29) in the Supplementary material, which gives the SNR of a realistic $N$ qubit scheme entirely in terms of the expectation and variance of the single qubit states |$\ket{0}$ and |$\ket{1}$ and the expectation and variance of the CNOT noise.
+The effects are summarised in Figs.
+2c and 2d.
+The SNR is calculated using (26) in the Supplementary Material, which gives the SNR of a realistic $N$ qubit scheme entirely in terms of the expectation and variance of the number of photons counted for an $N=1$ qubit scheme $K_{\ket{j},t}^{(1)}$ and the expectation and variance of the CNOT noise probability distribution.
 We see that we still obtain a significant space-time trade-off, allowing for significantly faster measurements, even in the presence of noise, demonstrating the robustness of this scheme.
+We note that our discussions have been framed in terms of the SNR, but we find essentially the same space-time trade-off if we consider the measurement infidelity: this is discussed in the Supplementary Material.
+In the perfect case this is shown analytically while for the imperfect case we present numerical results analogous to Fig.
+2c and 2d.
 There are other notable points worth highlighting.
@@
 3 in the Appendix in more detail.
-Similar to the perfect case, the rate of SNR increase improves with $N$, but it can now be a “super-improvement”; for small enough values of $N$ and CNOT noise $p$, the rate of speed up is better than (8) for the relevant time scales, that is the times before decoherence causes the measurement to deteriorate, as demonstrated by Fig.
+Similar to the perfect case, the rate of SNR increase improves with $N$, but it can now be a “super-improvement”; for small enough values of $N$ and CNOT failure rate $p$, the rate of speed up is better than linear (6) for the relevant time scales, that is the times before decoherence causes the measurement to deteriorate, as demonstrated by Fig.
 4 in the Appendix.
-V Conclusion
+VI Conclusion
 In this paper we have presented a general scheme for speeding up quantum measurement while maintaining quality.
@@
 We presented the main idea using ideal Poisson measurement statistics but emphasise that the scheme is much more general than this.
-In particular, by using characteristic noise models for the CNOTs and decoherence, the scheme was shown to be robust to imperfections while simultaneously able to achieve a “super-improvement” in speed-up, since it acts to suppresses decoherence, allowing for a higher achievable maximum quality.
+Moreover, by using characteristic noise models for the CNOTs and decoherence, the scheme was shown to be robust to imperfections while simultaneously able to achieve a “super-improvement” in speed-up, since it acts to suppresses decoherence, allowing for a higher achievable maximum quality.
 While we have illustrated the scheme with systems of qubits, it should be clear from the discussions above that the scheme is equally applicable to higher dimensional systems.
+Our case study was framed around state dependent fluorescence qubit measurements which are particularly appropriate to trapped-ion and cold-atom qubits.
+But, we would like to reiterate that our ideas are much more general.
+For example, our ideas also apply, when using multiple detectors, to superconducting qubits.
+We give a very brief discussion of this in the Supplementary Material.
+We also believe, just as in the earlier discussions, that we can embody our scheme for superconductors using a single detector; this is the subject of future work.
 Since speeding up quantum measurements is a highly relevant problem, we envisage that our scheme will have significant applications across quantum technologies.
-VI Note added
+VII Note added
 During preparation of this manuscript we became aware of the independent work of Petrosyan, Norrell, Poole and Saffman [31] which contains related ideas.
 Acknowledgements.
-VII Acknowledgements
+VIII Acknowledgements
 We acknowledge support from the EPSRC through the following grants: EP/T005386/1, EP/T001062/1, EP/Y005058/2, EP/Y004655/1, EP/L015730/1 and EP/R043957/1.
 We also acknowledge support from Agence Nationale de la Recherche (ANR-22-PNCQ-0002), the Royal Society (URF NFQI) and from a CIFAR Azrieli Global Scholarship.
-VIII Appendices
-VIII.1 Appendix on noise model
+IX Appendix on noise model
 We consider a noise model with two key ingredients: (i) Noise in the CNOT gates applied; (ii) decoherence of the individual systems. We outline here these two elements separately.
-For the CNOT noise, we consider a simple model inspired by neutral atom quantum computers.
+(i) For the CNOT noise, we consider a simple model inspired by neutral-atom quantum computers.
 In this model, whenever the system is in state |$\ket{1}$ and a CNOT is applied, there is a chance of failure, which causes the system to decay to |$\ket{0}$.
 More formally, we define a probability $p$, that immediately before a CNOT is implemented, |$\ket{1}$ will be sent to |$\ket{0}$.
-If this happens, the CNOT can not flip the state |$\ket{0}$ back to |$\ket{1}$.
+If this happens, the CNOT can not flip |$\ket{0}$ back to |$\ket{1}$.
 The noisy CNOT channel $\widetilde{C_{p}}$ acts on the basis states, where the first qubit is the control qubit, as
@@
 $\widetilde{C_{p}}(\ket{10}\bra{10})$ $=p\ket{00}\bra{00}+(1-p)\ket{11}\bra{11}.$ | . (10)
-Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{0}$ and |$\ket{1}$ of the system qubit.
+Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{\psi}=\ket{0}$ and |$\ket{\psi}=\ket{1}$ of the system qubit.
 Furthermore, we assume our ancillary qubits are initialised perfectly in |$\ket{0}$ so we are only interested in how the noisy CNOT gates will act on the states |$\ket{00}$ and |$\ket{10}$.
-Thus (VIII.1) defines the action on all states that occur in the scheme.
-When the system is in the state |$\ket{0}$, and we perform $N$ measurements, the initial state will be |$\ket{0}^{\otimes N}$.
+Thus (IX) defines the action on all states that occur in the scheme.
+When the system is in the state |$\ket{\psi}=\ket{0}$, and we have $N-1$ ancillary qubits, the initial state will be |$\ket{0}^{\otimes N}$.
 In this case, the action of $\widetilde{C_{p}}$ is to leave the state of all $N$ systems as |$\ket{0}^{\otimes N}$.
-When the system is in the state |$\ket{1}$, the initial state is |$\ket{1}\ket{0}^{\otimes(N-1)}$.
-Our scheme only depends upon the total number of photons emitted from each qubit (and does not depend upon which qubit emitted which photon).
-We thus only need to calculate the probability distribution for the total number of qubits in state |$\ket{1}$ after applying the $\widetilde{C_{p}}$ gates.
-There are numerous ways the gates can be compiled and different compilations will give different probability distributions for the number of qubits in state |$\ket{1}$.
-These different compilations are explored in more detail in the supplementary material.
+When the system is in the state |$\ket{\psi}=\ket{1}$, the initial state is |$\ket{1}\ket{0}^{\otimes(N-1)}$.
+Since our scheme only depends upon the total number of photons emitted from each qubit (and does not depend upon which qubit emitted which photon), we only need to calculate the probability distribution for the total number of qubits in state |$\ket{1}$ after applying the $\widetilde{C_{p}}$ gates.
+Figure 3: Increase in peak SNR from Fig.
+2c and 2d.
+This figure is extra detail for Figs.
+2c and 2d and uses the same parameters.
+We plot the peak SNR $S_{max}$ for $N=1$ to $N=10$ qubits.
+While more qubits results in a higher peak SNR, this trend flattens out for higher qubit numbers before eventually dropping.
+The slight unevenness in the plot is an odd/even effect of the cascade CNOT compilation, see Supplementary Material.
+Figure 4: Speed up from Fig.
+2c and 2d.
+This figure is extra detail for Figs.
+2c and 2d and uses the same parameters except for $p$, for which we plot two values.
+We plot the ratio $t^{(1)}_{S}/t^{(N)}_{S}$ for two different values of $p$, at $S=8$, for $N=1$ to $N=10$ qubits and the linear relationship $t^{(1)}_{S}/t^{(N)}_{S}=N$ (lime dashed line) achieved by the perfect Poisson distribution (6).
+The value $p=0.01$ (blue dots) is used in Figs.
+2c and 2d.
+The points are increasing, which shows we get a speed up, however they are below the dashed line so the speed-up is worse than the linear case.
+On the other hand, the values for $p=0.001$ (red crosses) are above the dashed line which demonstrates that, for CNOTs with a low enough failure rate, we can beat the decoherence of the qubit; it is possible to achieve a “super-improvement”.
+There are numerous ways to compile the gates and different compilations give different probability distributions for the number of qubits in state |$\ket{1}$, given an input state |$\ket{\psi}=\ket{1}$.
+These compilations are explored in more detail in the Supplementary Material.
 Let us denote by $\mathcal{T}^{(N)}_{p}(q)$ the probability distribution of the random variable $Q^{(N)}_{p}$ defining the number of qubits in state |$\ket{1}$ after applying the noisy CNOT gates.
@@
 Combining this with the probability distribution for $N$ qubits emitting $k$ photons (5), we find the resultant distribution using our scheme with noisy CNOT gates to be
-$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k)$
+$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k),$
 $\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(L_{\mu_{1}t}^{*q}*L_{\mu_{0}t}^{*N-q})(k)$
 $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)L_{(q\mu_{1}+(N-q)\mu_{0})t}(k).$ (11)
-For the decoherence noise, we consider a model whereby qubits in state |$\ket{1}$ probabilistically decay at a constant rate $\lambda$ to state |$\ket{0}$.
+(ii) For the decoherence, we consider a model whereby qubits in state |$\ket{1}$ probabilistically decay at a constant rate $\lambda$ to state |$\ket{0}$.
 After this happens, the qubit will emit photons at a different rate so the resultant distribution of photons will not be the same as a perfect Poisson.
 We are interested in the probability distribution for the number of photons emitted $k$ in time $t$.
-We model this behaviour by finding the probability a qubit in state |$\ket{1}$ will emit $k^{\prime}$ photons in time $t^{\prime}$ before decaying to |$\ket{0}$ and then emitting $k-k^{\prime}$ photons in time $t-t^{\prime}$, then integrating over $t^{\prime}$ and summing over $k^{\prime}$.
+We model this behaviour by finding the probability that a qubit in state |$\ket{1}$ will emit $k^{\prime}$ photons in time $t^{\prime}$ before decaying to |$\ket{0}$ and then emitting $k-k^{\prime}$ photons in time $t-t^{\prime}$, then integrating over $t^{\prime}$ and summing over $k^{\prime}$.
 We find the resultant probability distribution to be
@@
 The final integral can be readily evaluated numerically.
-Figure 3: Increase in peak SNR from Fig.
-2 (c) and (d).
-These Plots are generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1) with the cascade compilation of noisy CNOT gates defined in the Supplementary material.
-We use input parameters $\mu_{0}=3.5$, $\mu_{1}=14$, $\lambda=0.0041$ and $p=0.01$.
-We plot the peak SNR $S_{max}$ for $N=1$ to $N=10$ qubits.
-While more qubits result in a higher peak SNR, this trend flattens out for higher qubit numbers.
-Figure 4: Speed up from Fig.
-2 (c) and (d).
-Plot generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1) with the cascade compilation of noisy CNOT gates defined in the Supplementary material.
-We use input parameters $\mu_{0}=3.5$, $\mu_{1}=14$ and $\lambda=0.0041$.
-We plot the ratio $t^{(1)}_{S}/t^{(N)}_{S}$ for two different values of $p$, at $S=8$, for $N=1$ to $N=10$ qubits and the linear relationship $t^{(1)}_{S}/t^{(N)}_{S}=N$ (lime dashed line) achieved by the perfect Poisson distribution (8).
-The value $p=0.01$ (blue dots) is used in Fig.
-2 (c) and (d).
-The points are increasing, which shows we get a speed up, however they are below the dashed line so the speed-up is worse than the linear case.
-On the other hand, the values for $p=0.001$ (red crosses) are above the dashed line which demonstrates that, for CNOTs with a low enough failure rate, we can beat the decoherence of the qubit; it is possible to achieve a “super-improvement”.
-Our complete noise model consists of combining the noisy CNOT protocol (VIII.1) with the decaying Poisson distribution (15).
-Altogether, we we thus find that the probability distributions for the number of photons, using our scheme with $N$ systems measured, is given by
-$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k)$
+Our complete noise and decoherence model consists of combining the noisy CNOT protocol (IX) with the decaying Poisson distribution (12).
+Altogether, we find that the probability distributions for the number of photons emitted, using an $N$ qubit scheme, is given by
+$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k),$
 $\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(\mathcal{W}_{\mu_{1},\mu_{0},\lambda,t}^{*q}*L_{(N-q)\mu_{0}t})(k).$ (13)
-These are used to produce the results in Fig. 2 (c) and (d).
+These are used to produce the results in Figs. 2c and 2d where we use the cascade CNOT compilation, found in the Supplementary Material, for $\mathcal{T}^{(N)}_{p}(q)$.

### v3

@@
-Speeding up quantum measurement using space-time trade-off
+main_supplementary_material
 Abstract
@@
 This hardware-agnostic approach is broadly applicable to a range of quantum technology platforms and offers a route to accelerate mid-circuit measurement as required for effective quantum error correction.
-I Introduction
+.1 Introduction
 High-fidelity readout of individual quantum systems is an essential requirement for emerging quantum technologies including precision metrology and quantum computing.
@@
 In summary, our scheme is hardware agnostic and provides a robust, powerful and versatile method for speeding up quantum measurements using ancillary systems.
-II Case study: Poisson measurement statistics
+.2 Case study: Poisson measurement statistics
 In order to present the main idea, we will first focus on state dependent fluorescence qubit measurement which is the primary method used in trapped-ion [2, 24] and neutral-atom [1, 25, 26, 27] quantum computing.
@@
 Initially we consider $n=1$.
-In an idealised situation, the number of photons counted will follow a Poisson distribution
-$\mathcal{P}_{\ket{j},t}^{(1)}(k)=L_{\nu_{j}(t)}(k),$ (1)
-where $L_{\omega}(k)=\frac{\omega^{k}}{k!}e^{-\omega}$.
+In an idealised situation, the number of photons counted will follow a Poisson distribution ${\mathcal{P}_{\ket{j},t}^{(1)}(k)=L_{\nu_{j}(t)}(k)}$, where $L_{\omega}(k)=\frac{\omega^{k}}{k!}e^{-\omega}$.
 Recall the mean and variance of a Poisson distribution coincide and are both equal, $\langle K\rangle_{L_{\omega}}=\Delta K^{2}_{L_{\omega}}=\omega$.
 The two distributions, $\mathcal{P}_{\ket{0},t}^{(1)}$ and $\mathcal{P}_{\ket{1},t}^{(1)}$, will have an overlap which is the source of error in measurement quality when assigning outcomes; no threshold value can be defined which will give you a perfect measurement by correctly discriminating the two states.
-If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution,
-$L_{\nu_{j}(t_{1}+t_{2})}(k)$ $=\sum_{k=k_{1}+k_{2}}L_{\nu_{j}(t_{1})}(k_{1})L_{\nu_{j}(t_{2})}(k_{2}),$
-$=(L_{\nu_{j}(t_{1})}*L_{\nu_{j}(t_{2})})(k),$
-$=L_{\nu_{j}(t_{1})+\nu_{j}(t_{2})}(k),$
-i.e.
+If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution, ${L_{\nu_{j}(t_{1}+t_{2})}(k)=(L_{\nu_{j}(t_{1})}*L_{\nu_{j}(t_{2})})(k)=L_{\nu_{j}(t_{1})+\nu_{j}(t_{2})}(k)}$, i.e.
 the Poisson distribution convolves to itself.
@@
 This provides a quantification of the inverse trade-off between the speed and quality of a measurement.
-III Exact space-time trade-off in the idealised Poisson case
+.3 Exact space-time trade-off in the idealised Poisson case
 We now introduce a scheme which, rather than trading time for quality as above, allows us to trade time for “space” – by which we mean number of qubits.
@@
 We denote by $k_{1}$ and $k_{2}$ the number of photons detected from the first and second qubit detectors respectively and we input the state |$\ket{\psi}=\ket{j}$ into our scheme.
-The measurement statistics for the pair of outcomes $(k_{1},k_{2})$, given we are measuring the post-CNOT state |$\ket{jj}$ for a time $t$, are
-|$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2}).$ (4)
+The measurement statistics for the pair of outcomes $(k_{1},k_{2})$, given we are measuring the post-CNOT state |$\ket{jj}$ for a time $t$, are |${\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2})}$.
 The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, thus ${\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k)}$.
@@
 for all $S$, so the improvement in speed is exactly linear in qubit number. The speed up from our scheme is shown in Figs. 2a and 2b.
-IV General space-time trade-off
+.4 General space-time trade-off
 A key insight is that the scheme above applies much more generally than just to Poisson statistics arising from counting photons.
@@
 For this measurement we can calculate the SNR (1) between these two distributions.
-Let us assume the dependence with time, to leading order, is
-$S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}.$ (7)
-For the case of Poisson statistics above we saw that $\alpha~{}=~{}1/2$ however we may imagine an alternative dependence on time.
+Let us assume the dependence with time, to leading order, is ${S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}.}$ For the case of Poisson statistics above we saw that $\alpha~{}=~{}1/2$ however we may imagine an alternative dependence on time.
 Figure 2: Space-time trade-off for idealised and realistic devices.
@@
 In this case, a higher maximum distinguishability can be achieved by introducing more qubits and faster measurements can be made while also improving distinguishability.
-These plots are generated using the noise and decoherence model (IX), found in the Appendix.
+These plots are generated using the noise and decoherence model (I.1), found in the Appendix.
 We use input parameters $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate.
@@
 In the above we saw that, in an idealised version of the scheme involving perfect CNOTs, by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (3).
-Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, for perfect CNOTs we nevertheless have the fundamental property that the expectation and (due to independence) variance add,
-$\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}},$ $\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}.$
+Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, for perfect CNOTs we nevertheless have the fundamental property that the expectation and (due to independence) variance add, ${\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}}}$, ${\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}}$.
 It therefore follows that the SNR, using the scheme, will be
@@
 Thus, as long as $t^{\alpha}\sqrt{N}$ is constant, then the quality of the measurement is constant, and hence we still obtain a space-time trade-off.
-In particular, if $t_{S}^{(1)}$ is the duration of the measurement which achieves an SNR of $S$, then using the scheme, we can speed this up to a measurement time of $t^{(N)}_{S}$ given by
-$t_{S}^{(1)}/t^{(N)}_{S}=N^{1/(2\alpha)}.$ (9)
+In particular, if $t_{S}^{(1)}$ is the duration of the measurement which achieves an SNR of $S$, then using the scheme, we can speed this up to a measurement time of $t^{(N)}_{S}$ given by ${t_{S}^{(1)}/t^{(N)}_{S}=N^{1/(2\alpha)}}$.
 This recovers (4) in the case $\alpha=1/2$ but shows more generally, if $\alpha>1/2$, we will still achieve a trade-off between space and time using this scheme, albeit not as good as (4).
 Nevertheless, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (4).
-V Space-time trade-off in more realistic settings
+.5 Space-time trade-off in more realistic settings
 For our scheme to be useful in practical applications it is crucial to show that the speed-up is robust to realistic imperfection.
@@
 [29]).
-Additionally, we consider imperfections in qubit readout caused by qubits in state |$\ket{1}$ falling into |$\ket{0}$ during the measurement, such as due to depuming by off-resonant scattering in atomic fluorescence measurements [30].
+Additionally, we consider imperfections in qubit readout caused by qubits in state |$\ket{1}$ falling into |$\ket{0}$ during the measurement, such as due to depumping by off-resonant scattering in atomic fluorescence measurements [30].
 This error is included by applying a probabilistic decay at constant rate $\lambda$ during the measurement.
@@
 2c and 2d.
-The SNR is calculated using (26) in the Supplementary Material, which gives the SNR of a realistic $N$ qubit scheme entirely in terms of the expectation and variance of the number of photons counted for an $N=1$ qubit scheme $K_{\ket{j},t}^{(1)}$ and the expectation and variance of the CNOT noise probability distribution.
 We see that we still obtain a significant space-time trade-off, allowing for significantly faster measurements, even in the presence of noise, demonstrating the robustness of this scheme.
-We note that our discussions have been framed in terms of the SNR, but we find essentially the same space-time trade-off if we consider the measurement infidelity: this is discussed in the Supplementary Material.
+Fig.
+2 was generated using (I.2), in the Appendix, which gives a general expression for the SNR after using our scheme.
+It is applicable to essentially any arbitrary time-dependent noisy measurement process and accounts for arbitrary CNOT imperfections.
+We used (I.2) with the noise model outlined immediately above but, by using (I.2), the analogue of Fig.
+2 could be reproduced much more generally for any specific architecture.
+We note that our discussions have been framed in terms of the SNR, but we find essentially the same space-time trade-off if we consider the measurement infidelity: this is discussed in the Supplemental Material.
 In the perfect case this is shown analytically while for the imperfect case we present numerical results analogous to Fig.
@@
 4 in the Appendix.
-VI Conclusion
+.6 Conclusion
 In this paper we have presented a general scheme for speeding up quantum measurement while maintaining quality.
@@
 Our case study was framed around state dependent fluorescence qubit measurements which are particularly appropriate to trapped-ion and cold-atom qubits.
-But, we would like to reiterate that our ideas are much more general.
+We would like to reiterate that our ideas are much more general, as demonstrated by equation (I.2) in the Appendix.
 For example, our ideas also apply, when using multiple detectors, to superconducting qubits.
-We give a very brief discussion of this in the Supplementary Material.
+We give a very brief discussion of this in the Supplemental Material.
 We also believe, just as in the earlier discussions, that we can embody our scheme for superconductors using a single detector; this is the subject of future work.
+We have presented our results under the assumption that the CNOT duration is negligibly short compared to the measurement duration.
+The extent to which this is true is hardware specific, but it is straightforward to incorporate CNOT duration and, if it is sufficiently short, see that a trade-off will still be obtained.
+The relatively short two-qubit gate times compared to readout times in the following references demonstrate the advantage of our scheme to superconducting [31, 32] and neutral atom [1, 33] quantum computers while for trapped-ion quantum computers [34, 35] the advantage is less clear cut at present.
+We note however that this comparison is only relevant at the time of writing.
+Due to the fast development of quantum technologies the precise relation between two-qubit gate and readout durations is highly likely to change in the foreseeable future.
+Our view is that hardware providers will need to use every tool available to make their machines as effective as possible and our scheme has the potential to become a standard part of this toolbox.
+One of the most exciting applications of our scheme is to mid-circuit measurements, such as those used in error correction.
+As an important example, some form of parity measurement is key to many error correction schemes: here the parity of the qubits that is needed can be found by coupling those qubits to an ancilla using CNOTs, then measuring the ancilla.
+This ancilla measurement, and hence the whole error detection, can be sped-up by our scheme.
 Since speeding up quantum measurements is a highly relevant problem, we envisage that our scheme will have significant applications across quantum technologies.
-VII Note added
-During preparation of this manuscript we became aware of the independent work of Petrosyan, Norrell, Poole and Saffman [31] which contains related ideas.
+Note added.—-During preparation of this manuscript we became aware of the independent work of Petrosyan, Norrell, Poole and Saffman [36] which contains related ideas.
 Acknowledgements.
-VIII Acknowledgements
+I Acknowledgements
 We acknowledge support from the EPSRC through the following grants: EP/T005386/1, EP/T001062/1, EP/Y005058/2, EP/Y004655/1, EP/L015730/1 and EP/R043957/1.
 We also acknowledge support from Agence Nationale de la Recherche (ANR-22-PNCQ-0002), the Royal Society (URF NFQI) and from a CIFAR Azrieli Global Scholarship.
-IX Appendix on noise model
-We consider a noise model with two key ingredients: (i) Noise in the CNOT gates applied; (ii) decoherence of the individual systems. We outline here these two elements separately.
+I.1 Appendix A: On noise model
+We consider a noise model with two key ingredients: (i) Noise in the CNOT gates applied; (ii) decoherence of the individual systems.
+Initially we consider simple, albeit well-motivated models for both, but later in the Appendix we explain how the analysis can be extended to very general noise models.
+We first outline here these two ingredients separately in the simple model case.
 (i) For the CNOT noise, we consider a simple model inspired by neutral-atom quantum computers.
@@
 $\widetilde{C_{p}}(\ket{00}\bra{00})$ =|$=\ket{00}\bra{00}$ |
-$\widetilde{C_{p}}(\ket{10}\bra{10})$ $=p\ket{00}\bra{00}+(1-p)\ket{11}\bra{11}.$ | . (10)
+$\widetilde{C_{p}}(\ket{10}\bra{10})$ $=p\ket{00}\bra{00}+(1-p)\ket{11}\bra{11}.$ | . (A1)
 Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{\psi}=\ket{0}$ and |$\ket{\psi}=\ket{1}$ of the system qubit.
 Furthermore, we assume our ancillary qubits are initialised perfectly in |$\ket{0}$ so we are only interested in how the noisy CNOT gates will act on the states |$\ket{00}$ and |$\ket{10}$.
-Thus (IX) defines the action on all states that occur in the scheme.
+Thus (I.1) defines the action on all states that occur in the scheme.
 When the system is in the state |$\ket{\psi}=\ket{0}$, and we have $N-1$ ancillary qubits, the initial state will be |$\ket{0}^{\otimes N}$.
@@
 While more qubits results in a higher peak SNR, this trend flattens out for higher qubit numbers before eventually dropping.
-The slight unevenness in the plot is an odd/even effect of the cascade CNOT compilation, see Supplementary Material.
+The slight unevenness in the plot is an odd/even effect of the cascade CNOT compilation, see the Supplemental Material.
 Figure 4: Speed up from Fig.
@@
 There are numerous ways to compile the gates and different compilations give different probability distributions for the number of qubits in state |$\ket{1}$, given an input state |$\ket{\psi}=\ket{1}$.
-These compilations are explored in more detail in the Supplementary Material.
+These compilations are explored in more detail in the Supplemental Material.
 Let us denote by $\mathcal{T}^{(N)}_{p}(q)$ the probability distribution of the random variable $Q^{(N)}_{p}$ defining the number of qubits in state |$\ket{1}$ after applying the noisy CNOT gates.
@@
 $\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(L_{\mu_{1}t}^{*q}*L_{\mu_{0}t}^{*N-q})(k)$
-$=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)L_{(q\mu_{1}+(N-q)\mu_{0})t}(k).$ (11)
+$=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)L_{(q\mu_{1}+(N-q)\mu_{0})t}(k).$ (A2)
 (ii) For the decoherence, we consider a model whereby qubits in state |$\ket{1}$ probabilistically decay at a constant rate $\lambda$ to state |$\ket{0}$.
@@
 The final integral can be readily evaluated numerically.
-Our complete noise and decoherence model consists of combining the noisy CNOT protocol (IX) with the decaying Poisson distribution (12).
+Our complete noise and decoherence model consists of combining the noisy CNOT protocol (I.1) with the decaying Poisson distribution (A3).
 Altogether, we find that the probability distributions for the number of photons emitted, using an $N$ qubit scheme, is given by
 $\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k),$
-$\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(\mathcal{W}_{\mu_{1},\mu_{0},\lambda,t}^{*q}*L_{(N-q)\mu_{0}t})(k).$ (13)
-These are used to produce the results in Figs. 2c and 2d where we use the cascade CNOT compilation, found in the Supplementary Material, for $\mathcal{T}^{(N)}_{p}(q)$.
+$\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(\mathcal{W}_{\mu_{1},\mu_{0},\lambda,t}^{*q}*L_{(N-q)\mu_{0}t})(k).$ (A4)
+These are used to produce the results in Figs. 2c and 2d where we use the cascade CNOT compilation, found in the Supplemental Material, for $\mathcal{T}^{(N)}_{p}(q)$.
+I.2 Appendix B: On general space-time trade-off
+We now show how the scheme can be extended to very general noise/decoherence models with minimal assumptions.
+We allow for the CNOT gates to induce an arbitrary trace preserving quantum channel (i.e.
+we assume that the system does not leave the space spanned by the states |$\ket{0}$ and |$\ket{1}$ for each qubit) and we also assume that the measurement statistics are identical for each qubit.
+The SNR then depends only on (i) $P_{\ket{j},t}(k)$, the probability distribution of the random variable $K_{\ket{j},t}$ defining the measurement statistics of the basis state |$\ket{j}$ at time $t$ for an arbitrary time-dependent noisy measurement process and (ii) $T_{\ket{j}}^{(N)}(q)$, the probability distribution of the random variable $Q_{\ket{j}}^{(N)}$ defining the final number of qubits in state |$\ket{j}$ after applying the noisy CNOT gates to the input state |$\ket{j}$ and $N-1$ ancillary qubits initialised in state |$\ket{0}$.
+We note that the SNR will only depend on $T_{\ket{j}}^{(N)}(q)$, which is a limited statistic of the entangling CNOT process, so in order to compute the SNR we do not need full tomography of the quantum channel.
+In the Supplemental Material we show that the SNR is given by
+[…]
+×|$\times\absolutevalue{\langle Q\rangle_{\mathcal{T}_{\ket{0}}^{(N)}}+\langle Q\rangle_{\mathcal{T}_{\ket{1}}^{(N)}}-N}$ | (B1)
+where
+$\Delta K^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}=$ $\langle Q\rangle_{\mathcal{T}_{\ket{j}}^{(N)}}\Delta K^{2}_{P_{\ket{j},t}}$
+$+\left(N-\langle Q\rangle_{\mathcal{T}_{\ket{j}}^{(N)}}\right)\Delta K^{2}_{P_{\ket{\overline{j}},t}}$
+$+\left(\langle K\rangle_{P_{\ket{j},t}}-\langle K\rangle_{P_{\ket{\overline{j}},t}}\right)^{2}\Delta Q^{2}_{\mathcal{T}_{\ket{j}}^{(N)}}$
+and $\overline{j}=j+1~{}{\rm mod}~{}2$.
+Equation (I.2) is very general.
+Given the time dependent measurement statistics for a single qubit $P_{\ket{j},t}(k)$ and the CNOT noise as described by the probability distributions $\mathcal{T}_{\ket{j}}^{(N)}(q)$, it is possible to calculate the SNR for an $N$ qubit scheme at any time $t$.
+For example, experimentally we could collect the time dependent measurement statistics of a qubit in the state |$\ket{j}$ to find $P_{\ket{j},t}(k)$ and perform a - relatively limited - version of tomography on the CNOT gates to find the relevant noise distributions $T_{\ket{j}}^{(N)}(q)$.
+We could then use (I.2) to generate similar plots to Fig.
+2 which can be analysed to optimise over measurement duration, number of qubits and SNR.
+The full derivation and further discussion of (I.2) can be found in the Supplemental Material.